Solve the following word problem. Be sure to label the unknowns and to show the equation you use for the solution
Yuri has a board that is 98 in. long. He wishes to cut the board into two pieces so that one piece will be 10 in. longer than the other. What should the length of each piece be?

Answer by sarah_adam(201)

Length of the Board L = 98 in
But he wishes to cut the board in two pieces where one piece is 10 in long than the other
so let one piece be x in long
the other piece will be x + 10 in
so x + (x+10) = 98 in
2x = 98-10 = 88
x = 44 in
x + 10 = 54 in